OpenStudy (anonymous):

An apartment building has the following distribution of apartments: 1 bedroom 2 bedroom 3 bedroom 1st floor 3 1 1 2nd floor 2 2 2 3rd floor 1 4 1 If an apartment is selected at random, what is the probability that it is on the 2nd floor or has 2 bedrooms? A) 7/11 B) 6/11 C) 13/17 D) 11/17

OpenStudy (anonymous):

How many apartments have 2 bedrooms? Looking down the middle column, 1 + 2 + 4 = 7. How many apartments are on the second floor? Looking across the middle row, 2 + 2 + 2 = 6. But wait! We're counting something twice: two bedroom apartments on the second floor. Means you just need to subtract it (equivalent of only counting it once). 7 + 6 - 2 = 7 + 4 = 11. How many possibilities are there? How many apartments are there in this building? 3 + 1 + 1 + 2 + 2 + 2 + 1 + 4 + 1 = 17. 11 unique buildings match the criteria, and since there are 17 possibilities, the answer is D. 11/17.